OpenStudy (anonymous):

find the indicated probablity if p(a and b)=0.5 p(a)=0.7 and p(b)=0.5 find p(a or b)

OpenStudy (anonymous):

If p(a)=.7, then p=.7/a if p(b)=.5, then p=.5/a right so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you need to find p of a&b? so p(ab)?

OpenStudy (anonymous):

it only says p( a or b)

OpenStudy (m):

are you sure p(a and b) = 0.5

OpenStudy (anonymous):

yeah that's what it says if P(a and B)=0.5

OpenStudy (m):

do you have the answer? is it 0.7?

OpenStudy (anonymous):

I am trying to figure it out still

OpenStudy (anonymous):

I think for 'or' you add the probibility

OpenStudy (anonymous):

would it be 0.7?

OpenStudy (m):

i think so

OpenStudy (anonymous):

.5/a+.7/b then? So p=.5/a and p=.7/b Would that be the final answer? Trying to understand this one too.

OpenStudy (m):

if A = .5 and if B = .7 and P(A and B) = .5 that means A overlaps completely over B P(A or B) = P(A) + P(B) + P(A and B) = 0.7

OpenStudy (dumbcow):

M is correct

OpenStudy (anonymous):

Ignore my answer then, not bad for a first shot at a probability problem though ;)

OpenStudy (dumbcow):

wait typo should be a "-" P(A or B) = P(A) + P(B) - P(A and B) = 0.7

OpenStudy (m):

last equation i wrote is wrong sorry

OpenStudy (anonymous):

Ok I have a new single die one if u guys wanna look at it